# Multi-Object Tracking with Camera-LiDAR Fusion for Autonomous Driving

### v1

@@ III Experimental results
  • Two Robosense M1 solid state LiDARs, each one guaranteeing an horizontal Field Of View (FOV) of ${}^{\circ}$ and a vertical one of ${}^{\circ}$, with a resolution of ${}^{\circ}$. The two LiDARs are positioned with ${}^{\circ}$ of overlap, in order to obtain a total horizontal FOV of ${}^{\circ}$. The two LiDARs are synchronized using the gPTP protocol, allowing for the merging of sensor outputs before the processing step.
  • A Zed2i camera, with an horizontal FOV of ${}^{\circ}$ and a focal length of 2.1mm.
@@ III-A KITTI MOT benchmark
Class
HOTA (%)
DetA (%)
AssA (%)
LocA (%)
MOTA (%)
Car
76.784
76.378
77.27
87.498
88.472
Pedestrian
44.737
47.019
42.813
79.536
43.928
@@ III-B State estimation accuracy
pos [$m$] $\psi$ [${}^{\circ}$] $v$ [$m/s$] $\omega$ [${}^{\circ}/s$] pos [$m$] $\psi$ [${}^{\circ}$] $v$ [$m/s$] $\omega$ [${}^{\circ}/s$] pos [$m$] $\psi$ [${}^{\circ}$] $v$ [$m/s$] $\omega$ [${}^{\circ}/s$]
@@
Single-modal (camera)
Single-modal (LiDAR)
Multi-modal       (camera+LiDAR)
RMSE pos [$m$]
0.1804
0.1671
0.1378
$\psi$ [${}^{\circ}$]
2.171
3.035
2.329
$v$ [$m/s$]
0.2782
0.2811
0.273
$\omega$ [${}^{\circ}/s$]
3.674
4.517
4.15
MAE pos [$m$]
0.1271
0.097
0.085
$\psi$ [$\circ$]
1.663
1.877
1.786
$v$ [$m/s$]
0.2016
0.201
0.196
$\omega$ [${}^{\circ}/s$]
2.809
3.286
3.185
MaAE pos [$m$]
1.015
0.845
0.797
$\psi$ [${}^{\circ}$]
11.017
31.939
9.614
$v$ [$m/s$]
0.8787
1.238
0.873
$\omega$ [${}^{\circ}/s$]
16.227
22.803
14.744

### v2

@@ III Experimental results
  • Two Robosense M1 solid state LiDARs, each one guaranteeing an horizontal Field Of View (FOV) of 120 and a vertical one of 25, with a resolution of 0.2. The two LiDARs are positioned with 30 of overlap, in order to obtain a total horizontal FOV of 180. The two LiDARs are synchronized using the gPTP protocol, allowing for the merging of sensor outputs before the processing step.
  • A Zed2i camera, with an horizontal FOV of 120 and a focal length of 2.1mm.
@@ III-A KITTI MOT benchmark
Class HOTA (%) DetA (%) AssA (%) LocA (%) MOTA (%)
Car 76.784 76.378 77.27 87.498 88.472
Pedestrian 44.737 47.019 42.813 79.536 43.928
@@ III-B State estimation accuracy
pos [$m$] $\psi$ [] $v$ [$m/s$] $\omega$ [${}^{\circ}/s$] pos [$m$] $\psi$ [] $v$ [$m/s$] $\omega$ [${}^{\circ}/s$] pos [$m$] $\psi$ [] $v$ [$m/s$] $\omega$ [${}^{\circ}/s$]
@@
Single-modal (camera) Single-modal (LiDAR) Multi-modal       (camera+LiDAR)
RMSE pos [$m$] 0.1804 0.1671 0.1378
$\psi$ [] 2.171 3.035 2.329
$v$ [$m/s$] 0.2782 0.2811 0.273
$\omega$ [${}^{\circ}/s$] 3.674 4.517 4.15
MAE pos [$m$] 0.1271 0.097 0.085
$\psi$ [$\circ$] 1.663 1.877 1.786
$v$ [$m/s$] 0.2016 0.201 0.196
$\omega$ [${}^{\circ}/s$] 2.809 3.286 3.185
MaAE pos [$m$] 1.015 0.845 0.797
$\psi$ [] 11.017 31.939 9.614
$v$ [$m/s$] 0.8787 1.238 0.873
$\omega$ [${}^{\circ}/s$] 16.227 22.803 14.744